OpenStudy (anonymous):

How do you find the vertex form of f(x)=-2x^2-6x+10

OpenStudy (tkhunny):

Completing the square! Go!

OpenStudy (anonymous):

thats what im trying to do but I dont understand how

OpenStudy (tkhunny):

Two hints. 1) Think long and hard on this: \((a+b)^{2} = a^{2} + 2ab + b^{2}\) 2) f(x)=-2x^2 - 6x + 10 = -2(x^2 + 3x + ______) + 10 - (-2)(______) Just fill in that big blank with the same value. What is the value? Please see hint number 1).

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

still not really fully understanding

OpenStudy (tkhunny):

Hint #3) You have 2b and you need b^2. Give it another go.

OpenStudy (anonymous):

I'm sorry no im not getting the second number that is needed.

OpenStudy (anonymous):

i got -2(x-__)^2

OpenStudy (tkhunny):

One more. 2b = 3 Find b^2

OpenStudy (anonymous):

2.25

OpenStudy (anonymous):

sorry for taking so much of your time

OpenStudy (tkhunny):

Don't use 2.25. Stick with fractions. b = 3/2 b^2 = 9/4 f(x)=-2x^2 - 6x + 10 = -2(x^2 + 3x + 9/4) + 10 - (-2)(9/4) Can you simplify that right-hand expression?

OpenStudy (anonymous):

didnt notice the message earlier i sticked with the decimal and got -2(x^2+1.5)+14.5

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

-2(X+9/4) + 7

OpenStudy (tkhunny):

(x^2 + 3x + 9/4) = (x + 3/2)^2 10 + (2)(9/4) = 10 + 9/2 = 20/2 + 9/2 = 29/2 Work on your algebra. You'll get it.